OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#5763 — Optical switching buckle
Scheduled Maintenance Report for Web Cloud
Completed
We are going to launch the first optical buckle tonight around 11h00. that's why we have to cut the NRA, PIL, BLE, BOI, FAC, SEC and LIB.

This launching is expected during an important part of night when internet access are cut.

Thank you for your understanding !

We will keep you informed of all different steps.


Update(s):

Date: 2011-09-09 16:52:21 UTC
The problem is at the level of the routing.

We are changing the router. The wiring is in progress.
All the NRA will be recables in the following minutes.

Date: 2011-09-09 16:49:54 UTC
There was latency periods and packets losses on the NRA of BLE59.
We intervened on the spot to change certain passive euipments on the site.

We have a better one but we still have some problems.

We are still looking at the level of the routing.

Date: 2011-09-09 05:51:17 UTC
NRA, LIB, SEC, FAC, BOI, BLE and PIL Routing are again UP.


Date: 2011-09-09 05:50:21 UTC
All optical equipments are implemented. We need a link to finalise the buckle. We faced a problem while implementation. All sites are fixed and calibrated. We will accomplish the routing and work tonight.


Date: 2011-09-09 05:47:53 UTC
SEC FAC and BOI sites are related optically. We are setting the BLE site in order to include it in the optical buckle. We are re-establishing meanwhile the traffic.


Date: 2011-09-08 23:44:20 UTC
PIL and LIB sites are attached optically on the POP of Roubaix.


Date: 2011-09-08 23:43:35 UTC
Optical fibers are switched on the optical chassis. We are starting work to mount the optical rings.
Posted Sep 08, 2011 - 23:42 UTC